Consider the following partial analysis of variance AN OV A table.

Source of variation d f SS M S F

Treatments    75 25   

Error    60   

TOTAL 19 135

The test statistic is

1. 28:75

2. 6:67

3. 0:15

4. 100

5. 6:067


1
Expert's answer
2021-09-02T23:53:29-0400

Df treatments = SStreatments/MStreatments = 75/25 = 3

Df error = DFtotal- Dftreatments = 19 - 3 = 16

MSerror= SSerror/Dferror = 60/16 = 3.75

Test statistic F = MSTreatments /MSError = 25/3.75 = 6.67

Answer = 6.67